OpenStudy (anonymous):

4.2x - 3 = .5 (8.4x + 6) and 2/3x + 1 = 2/3x -1 Just lead me in the right direction please mates? I don't need the answer, just to know how to do these.

Elsa213 (elsa213):

\(\Large\frak\color{Magenta}{\heartsuit Welcome~To~Openstudy\heartsuit}\) \(\color{Purple}{1.~~~Please~read~the~Code~Of~Conduct~when~You~have~a~chance}\) http://openstudy.com/code-of-conduct \(\color{pink}{T}\color{Green}{^h}\color{Orange}{\heartsuit ^a}\color{Red}{n}\color{Purple}{^K}\color{lime}{~~^Y}\color{yellow}{o}\color{blue}{^u}\)

OpenStudy (amistre64):

is it 2 different questions? or a system of equations?

OpenStudy (anonymous):

two different

OpenStudy (amistre64):

you have 3 basic algebra properties to play with: identites, inverses, and distribution/factoring

OpenStudy (amistre64):

4.2x - 3 = .5 (8.4x + 6) distribution: a(b+c) = ab+ac factoring: ab+ac = a(b+c) one idea is to distribute to start with. what do we get as a result?

OpenStudy (anonymous):

Well, I'm not very good a t math. But the way I was told was to do something like 4.2x - 3 = .5(8.4x + 6) 4.2x - 3 = 4.2x + 3 -4.2x -3 = -4.2x +3 And this is where I get lost cause I either get -6 or +6

OpenStudy (anonymous):

Same goes for the second, I get either -2 or +2

OpenStudy (anonymous):

Neither make sense when I try to check.

OpenStudy (amistre64):

.2x - 3 = 4.2x + 3 is fine to play with lets start getting the x parts all by themselves. -3 + ? = 0

OpenStudy (anonymous):

+3?

OpenStudy (amistre64):

teh 4 didnt copy over, but yes .... so lets add 3 to each side 4.2x - 3 = 4.2x + 3 +3 +3 ----------------- 4.2x = 4.2x + 6 is there any number that you can think of that is 6 more than itself?

OpenStudy (anonymous):

Wait, but aren't we supposed to do subtract x from both side? I know I'm supposed to get x all by itself, but isn't that when there are two variables, I have to subtract/add from one?

OpenStudy (amistre64):

you can if you need to, but math is logical

OpenStudy (amistre64):

if we subtract 4.2x from each side we get: 0 = 6 which is not a logical construction

OpenStudy (amistre64):

this means that there is no solution that will make it true

OpenStudy (anonymous):

Mmm, I guess that makes sense.

OpenStudy (amistre64):

2/3x + 1 = 2/3x -1 is x on the bottom or beside the 2/3 ?

OpenStudy (amistre64):

either way, this is the same logic. 2/3x is just some number add 1 to each side ... is there a number that is equal to 2 more than itself?

OpenStudy (amistre64):

or rather, when does 0=2?

OpenStudy (anonymous):

0 never equals something. Wait, hold on, I'm trying to figure out your smartness first. Lemme try answering the first one with this new info. I'll give you what I got. gimem about a 2 minutes.

OpenStudy (amistre64):

if my internet holds up, ill be around :)

OpenStudy (anonymous):

okay so, I'm still stuck on the first question. 4.2x = 4.2x + 3 Am I on the right track?

OpenStudy (amistre64):

+6 not +3 but yeah

OpenStudy (anonymous):

then I subtract X?

OpenStudy (amistre64):

we have 4.2x on each side, to keep things equal, what we do to one side we do to the other. if we remove 4.2x from one side, we remove it from the other. subtracting x is doable, but doesnt really lead us anyplace useful

OpenStudy (amistre64):

4.2x = 4.2x + 3 4.2x-x = 4.2x-x + 3 3.2 = 3.2x + 3

OpenStudy (amistre64):

i spose it gets rid of 1x ... 3.2 more to go :)

OpenStudy (anonymous):

So, then. I can't subtract x, cause then I won't have an x.

OpenStudy (amistre64):

and i trust your smart enough to see thru the typos

OpenStudy (amistre64):

you can subtract up to 4.2 xs after that there are no more that can be subtracted

OpenStudy (amistre64):

4.2x = 4.2x + 6 -4.2x -4.2x --------------- 0x = 6 so we get to 0=6

OpenStudy (anonymous):

SO, I was sorta right? 6 is the answer?

OpenStudy (anonymous):

But when I try to check, substitution for 6, I get different answers.

OpenStudy (amistre64):

6 isnt the answer, the answer is, what value of x makes: 0x = 6 true?

OpenStudy (amistre64):

so, 0, times what, is equal to 6?

OpenStudy (anonymous):

Argh! But 0 times anything is always 0.

OpenStudy (anonymous):

I dun geeeet it. Should I go in for help period?

OpenStudy (amistre64):

a math sentence can be as gibberish and meaningless as any other sentence. when a math sentence is meaningless, or nonsensical, we conclude that there is no answer that will suffice.

OpenStudy (amistre64):

0x = 6 has no solution.

OpenStudy (amistre64):

therefore: 4.2x - 3 = .5 (8.4x + 6) has no solution and is a nonsensical math sentence

OpenStudy (amistre64):

geometrically, this represents 2 lines that never meet

OpenStudy (anonymous):

Wait, is there an answer to the equations though?

OpenStudy (amistre64):

the answer is that there is no solution to the equation

OpenStudy (anonymous):

That's new. Kay thanks